OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 2x + 3y = 3 3x – 2y = 11 What is the solution to the system? A. (–3, 3) B. (–1, –7) C. (1, –4) D. (3, –1)

OpenStudy (anonymous):

Do you have a graphing calculator?

OpenStudy (anonymous):

No /:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

plug in the answers into both equations...2(-3)+3(3)=3 and the second one also 3(-3)-2(3)=11

OpenStudy (anonymous):

-6+9=3 and -9-6=11

OpenStudy (anonymous):

yes, how about the others?

OpenStudy (anonymous):

(-1,-7) dont work -3+14=11 but -2-21 doesnt =3

OpenStudy (anonymous):

Oh I thought you just wanted me to do the ones you gave me ok hold on lemme do them real quick

OpenStudy (anonymous):

When in doubt.....I work em out :O)

OpenStudy (anonymous):

lol letter D works for both equations

OpenStudy (anonymous):

Im graphing them now

OpenStudy (anonymous):

the solution is where the two lines cross

OpenStudy (anonymous):

Ok thank you so much! :)